OpenStudy (anonymous):

Using the recursive formula, f(n+1)=2*f(n)+1, and given that f(1)=3, calculate the product of the digits in the number f(17).

OpenStudy (amistre64):

what is f(2)?

OpenStudy (amistre64):

\[f(1+1)=2~f(1)+1\] \[f(2)=2(3)+1\] knowing f(2), what is f(3)? and so on until you get to f(17)

OpenStudy (loser66):

I think you have to solve for f(n). You have the RHS = constant, that means you have to look for a f(Pn) = A and then 2f(Pn) =2A -------------------- f(Pn)) - 2 f(Pn) =-A =1 ---> A = -1 LHS you have the form of f(n+1) -2f (n) = 0 .The characteristic equation for that is L(lamda) - 2 = 0 ---> L =2 So the form of the function f(n) for the LHS is C *2^n Now, combine both sides, f(n) =( C * 2^n) -1 when you have f(1) =3 replace to get C (C*2) -1 =3 ---> C = 2 Replace to form of f(n). You get the form f(n) = (2 * 2^n )-1 That is the final form to calculate the f of whatever value of n, It no longer depends on the previous term. if you want to calculate f(17)? f(17) = (2 * 2^17) -1= 262143 Hope this what you need

OpenStudy (loser66):

@amistre64

OpenStudy (amistre64):

ida put it in usual notation:\[a_n=2a_{n-1}+1~:~a_1=3\] \[a_2=2(3)+1=7\] \[a_3=2(7)+1=15\] \[a_4=2(15)+1=31\] \[a_n=2(31)+1=63\] 3 7 15 31 63 4 8 16 32 \[a_n=2^n-1\]

OpenStudy (amistre64):

might have my index alittle outta sight

OpenStudy (amistre64):

\[a_n=2^{n+1}-1\]might be better for me

OpenStudy (amistre64):

which is what you have yes

OpenStudy (loser66):

you miss something, since a(n) = a(n -1) +1 so a (n+2) = a(n+1) +1 = a(n) +1 +1

OpenStudy (loser66):

thnks for medal.

OpenStudy (amistre64):

ive seen your method about, but i have never had the time to actually learn it :)

OpenStudy (loser66):

I don't know, from my experience, I have to put everything in neat logic

OpenStudy (loser66):

so, you are genius in the field you never learn before, hat off

OpenStudy (amistre64):

my teachers love me ;)

OpenStudy (loser66):

how lucky you are, my profs hate me most